OpenStudy (anonymous):

help please :) solve: 11/5.9 = 15/x

OpenStudy (anonymous):

wouldn't you just cross multiply because your solving for x?

OpenStudy (anonymous):

im not sure lol @jennyk625

OpenStudy (anonymous):

I believe you just take 5.9 times 15 and than divide by 11 and than you get x solved for.

OpenStudy (phi):

there are a few ways to solve this. It would be good to learn one of them if you cross multiply (multiply the bottom of one side times the top of the other side) you get 11x= 15*5.9 now divide both sides by 11 to get x * 11/11 = 15*5.9/11 or x= 15*5.9/11

OpenStudy (anonymous):

its 8? @phi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thankyou !! @jennyk625 @phi